OpenStudy (anonymous):

If (x+2)/2 = y/5, then which of the following must be true? x/2 = (y-2)/5 x/2 = (y-5)/5 (x+2)/5 = y/25

OpenStudy (lgbasallote):

i cannot see where those three equations came from o.O but im sure @Diyadiya does

OpenStudy (anonymous):

the problems on the bottom answer choices A B or C

OpenStudy (lgbasallote):

i know...and i dont know where they came from

OpenStudy (anonymous):

Well thats all the ? ask?

OpenStudy (lgbasallote):

oh...nevermind...found it :p

OpenStudy (anonymous):

oh ok

OpenStudy (lgbasallote):

cross multiply \[\frac{(x+2)}{2} = \frac{y}{5}\] first what do you have?

OpenStudy (anonymous):

ok! then?